Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

National Squad Training

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Comber Rifle Club, Comber, Co.Down
Date June 23-24, 2001
Discipline 50m Three Position Rifle, 50m Prone Rifle, 10m Air Rifle

DURC Attendees

Mark Dennehy Steven Watterson

John Keeney

Report

Three DURC members, (Club Captain Steven Watterson, Ammunition&Equipment Officer John Keeney and Club Secretary Mark Dennehy) attended the Irish National Squad Training weekend in Comber, Co.Down this weekend. This weekend proved somewhat different in format from the normal weekends as the location meant that for most shooters it became a residential course with the DURC and Wilkinstown shooters staying on the Comber range for the weekend. Several hours a day were spent training with the National Coach Barry Dagger supervising and instructing. It was felt that this was an excellent weekend by those that went, and many lessons about 50m outdoor shooting were learnt, culminating in John Keeney shooting a perfect score – the first time he has done so in a 50m range.

The scheduling of the next National Squad training session has not yet been decided, but when arranged it is anticipated that DURC will once again be attending, especially since Mark is currently on the National Squad.

Misé le meas,

Mark Dennehy, DURC Secretary